A balloon has an internal pressure 1.014 atmospheres and a radius 6.7 centimetres. What is the resultant pressure force on the elastic membrane


1
Expert's answer
2021-09-17T09:22:49-0400

Given:

"P=\rm 1.014\: atm=102 700\: Pa"

"R=\rm 6.7\: cm=0.067\: m"


The pressure force on the elastic membrane is given by

"F=P*A=P*4\pi R^2""F=\rm 102 700\: Pa*4*3.14*(0.067\: m)^2=5790\: N"

But, due symmetry, the resultant pressure force on the membrane is equal to zero.

Answer: "F=0\:\rm N"